# Predicting Field Experiments with Large Language Models

### v1

@@ Abstract
Large language models (LLMs) have demonstrated unprecedented emergent capabilities, including content generation, translation, and the simulation of human behavior. Field experiments, despite their high cost, are widely employed in economics and the social sciences to study real-world human behavior through carefully designed manipulations and treatments. However, whether and how these models can be utilized to predict outcomes of field experiments remains unclear. In this paper, we propose and evaluate an automated LLM-based framework that produces predictions of field experiment outcomes. Applying this framework to 319 experiments drawn from renowned economics literature yields a notable prediction accuracy of 78%. Interestingly, we find that performance is highly skewed. We attribute this skewness to several factors, including gender differences, ethnicity, and social norms.
1 Introduction
@@
Field experiments allow researchers to manipulate variables of interest in a real-world setting, establishing causal relationships between interventions and outcomes. They typically begin by designing an intervention aligned with a specific research question, randomly assigning participants to treatment or control groups, and then measuring outcomes under natural conditions. The resulting data are collected and analyzed to evaluate the causal impact of the intervention Levitt and List (2009). It is adopted by a wide range of disciplines across academia and industry such as human behavior understanding Kao et al. (2024), marketing strategies making Hsu and Karahalios (2024), impact of machine learning on human Reinmund et al. (2024).
@@
With the advancement of Generative AI, many scholars believe it is possible to leverage LLMs’ content generation and reasoning ability to simulate human responses and behavior Taubenfeld et al. (2024). Existing papers have confirmed the alignments between simulated data generated by LLMs and real data collected from human participants across various aspects, including human responses, traits Dubois et al. (2024), moral standards Dillion et al. (2023), preferences Namikoshi et al. (2024), and emotions Gao et al. (2023). Based on those encouraging results, scholars have successfully steered LLMs to replicate existing lab experiments across several disciplines, including psychology Aher et al. (2023), sociology Leng and Yuan (2023); Manning et al. (2024), and economics Aher et al. (2023); Horton (2023), which aim to replicate existing lab experiments by treating LLMs as participants of lab experiments.
@@
However, these works have several disadvantages. First, these works only focused on lab experiment settings. But field experiments are inherently more challenging to conduct than laboratory experiments, due to diverse participant backgrounds, complex workflows, and multifaceted treatment designs. Second, previous studies mainly replied on manual process and only showcased a handleful experiments most related to social psychology. In order to examine LLMs’ capabilities robustness and generalizability in experiment simulation, we need to test it against a much larger scale and broader range of experiments.
In this paper, we fulfill the literature gap by proposing an automated LLM framework based on GPT to directly predict the outcome of field experiments. This framework has several components: an information extraction module, a variant generation module, and a prediction module. We test this framework on over 300 field experiments containing a total of 1638 conclusions. Using an LLM-powered multilevel filtering mechanism, we get these experiments screening from over 6,500 papers related to “field experiment” published in premier academic journals from 2000 to 2024.
Without any alignment techniques or fine-tuning, our framework achieves an average prediction accuracy of 78%. Specifically, for each conclusion, the framework generates two false variants, then inserts the true conclusion and two distractors into a prediction prompt template with a Chain-of-Thought design, prompting the LLM to select one of the three options. Interestingly, we also find that the prediction results are highly skewed. For example, our framework achieves nearly 100% prediction accuracy on 73% of conclusions while it completely fails to predict 16% of the conclusions with close to 0% accuracy. Further investigation on that skewness reveals that the LLM-based framework has limitations on predicting experiment with a context of gender differences, ethnicity, and social norms.
@@
Our research makes several contributions. First, we extend the literature on LLMs’ emergent capabilities by demonstrating that LLMs can simulate field experiments by predicting conclusions. While using LLMs to simulate human behavior is well studied, to the best of our knowledge, this is the first work in literature that considers simulating field experiments that require a more complex environment setting and workflow design.
@@
Second, our proposed framework allows the prediction of field experimental conclusions in a fully automated, large-scale fashion. As a result, our framework’s capability is more convincing and can be generalized to a wider range of applications.
Third, by comparing the predictions and ground truth from existing field experiments, this paper identifies infeasible scenarios where the LLM underperforms in prediction tasks, revealing the limitations of the LLM in field experimental simulations caused by confound reasons.
2 Literature Review
@@
While past literature mainly focused on agent-based social simulation Leung et al. (2024); Tennant et al. (2023); Zehmakan et al. (2023), there is an increasing trend to adopt LLMs as simulation tools. Existing studies have found that LLMs’ ability to simulate human behavior source from their possession of human-like reasoning skills and their adaptivity to personas of diverse characters Shanahan (2024); Taubenfeld et al. (2024). Upon those features of LLMs, Aher et al. Aher et al. (2023) proposed a concept of Turing Experiment that LLMs can be profiled as synthetic participants of experiments with integrated prompt of experimental settings and demographic information. Similarly, Horton Horton (2023) proved LLMs’ ability to simulate lab experiments and promoted it as a method of experimental pilot testing. The usage of LLMs in Horton Horton (2023)’s work is similar to Aher et al. Aher et al. (2023), consisting of two stages: prompting LLMs as synthetic experimental participants and collecting responses from conversations, reporting a few successful simulation cases of lab experiments. Leng & Yuan Leng and Yuan (2023) harnessed a three-phase procedure to complete the lab experimental simulation, which includes the initialization phase, interaction phase, and decision analysis phase. In the initialization phase, separate conversations of GPT-4 are prompted as vanilla experimental participants without specifying demographic information. Then, in the interaction stage, synthetic participants are prompted with the actions of other participants and asked to take actions and rationales according to the experiment design. Last, agents’ actions and rationales are collected for analysis to conclude. Leveraging this procedure, the study simulates five existing lab experiments. Manning et al. Manning et al. (2024) proposed an automated framework for various lab experimental simulations. Although the entire workflow is divided into seven steps, it is essentially similar to the manual procedures of the three papers above. However, the sole input is the backstory of the experiment. Based on the input information and powered by an LLM, the framework continues with subsequential steps: identifying dependent and independent variables, generating treatment values, profiling agents, organizing the interaction of agents, collecting data, and establishing causal relationships and conclusions. The paper reports the results of four social scenarios.
@@
In summary, while existing studies discussed above pioneered the application of Large Language Models (LLMs) to laboratory experiment simulations, several research gaps remain unaddressed. For instance, existing methods are not tested on field experiments that are inherently more complex than laboratory experiments, encompassing diverse participant backgrounds, more intricate workflows, and multifaceted treatment designs. Besides, none of these studies examined the impact of bias on simulation outcomes, despite extensive discussions of AI bias at IJCAI Wang et al. (2023); Du and Zhang (2024). Given that recognized biases in LLMs—such as gender Khorramrouz et al. (2023) and social norm bias Ren et al. (2024)—can compromise performance on downstream tasks Harris et al. (2022), it is crucial to identify scenarios in which bias undermines the fidelity of experimental simulations. Another limitation is that these studies tested their simulation strategies on only a small number of experiments, restricting generalizability. To address these gaps, we evaluated our proposed framework at scale on field experiments from published papers, demonstrating not only that LLMs can accurately predict the outcomes of established experiments, but also clarifying boundary conditions under which they cannot provide reliable experimental predictions. Furthermore, the only existing automated simulation framework Manning et al. (2024) is incompatible with scenarios involving complex human-object interactions or precise experimental control. By contrast, our framework offers a broadly adaptable approach that supports experiments across a wide range of contexts.
3 Data Collection and Filtering
@@
To test LLMs’ ability to predict field experiments in a large-scale fashion, we collected existing field experiments from papers published in the top 4 journals in economics, such as The Review of Economic Studies, American Economic Review, Journal of Political Economy, and The Quarterly Journal of Economics. It is worth noting that we focus on field experiments in economics because they are generally with a larger scale and reflect more generic human behaviors compared to field experiments in other disciplines Levitt and List (2009).
Figure 1: The Data Collection Workflow
@@
Figure 1 shows the data collection and filtering workflow. Initially, 6544 papers containing keywords related to field experiments published between 2000 and 2024 were selected. Then, we applied a two-layer verification process powered by Claude (Claude-3-opus-20240229). First, we prompted the title and abstract of each paper to Claude and asked it to judge if the paper designs and implements a field experiment. Upon this, the second verification prompted the entire paper to Claude, asking the same question. This strategy balances the accuracy of verification and the cost of calling Claude, as prompting the entire paper exponentially increases the cost. It is also worth mentioning that using Claude as the verification tool, instead of GPT, prevents potential data leakage as GPT is leveraged as the prediction tool in our framework. The distribution of the selected papers is shown in Figure 2.
Figure 2: A Summary of Qualified Papers
4 Framework
@@
We present the details of our automated framework for predicting field experimental conclusions in this section. Figure 3 shows the workflow. Overall, our framework is divided into three stages: Extraction from Papers, Variant Generation, and Prediction. Notably, Claude (Claude-3-opus-20240229) powers all preprocess tasks in the first two stages, whereas GPT completes the prediction at the last stage. We use two different LLMs in different stages to prevent potential data leakage from one another.
Figure 3: Prediction Framework
4.1 Extraction
Figure 4: Prompt for Extraction
@@ 4.1 Extraction
Specicifically, the framework uses Claude to extract information related to a field experiment from the selected paper. To realize this, the framework leverages a manually crafted prefixed prompt, which has proven to be efficient for various downstream tasks Brown et al. (2020). As shown in Figure 4, the prompt template contains a placeholder “Paper”, an information form consisting of bullet points from “A” to “G”, and clear instructions that ask the LLM to extract information according to the form from the paper. As underscored in the prompt, the first six bullet points “A” to “F” are key experimental settings that shape the experiment context, whereas the last point “G” is about conclusions that are true outcomes in the prediction task.
@@
Based on the response from Claude, the framework formulates “Experiment Settings” directly from bullet points “A” to “F”, while it polishes point “G” to generate “Experiment Conclusions”. Specifically, raw response regarding point “G” is a paragraph containing multiple conclusions of the field experiment. To separate that paragraph into standalone conclusions, the framework calls a new Claude session, prompting raw paragraph and related instructions to it, finally getting “Experiment Conclusions” in return. Breaking complex tasks into subtasks improves the performance of LLM-driven workflows Zhang and Gao (2023), which is the main reason for completing extraction and separation of conclusions in different Claude sessions.
Furthermore, we compared the extracted information by our approach, GPT-4, and Elicit (https://elicit.com/) over the same paper Bott et al. (2020). Elicit is an LLM fine-tuned on scholarly databases that can comprehend papers and extract related information Bail (2024), which could be a benchmark. As a result, whereas GPT-4 and Claude-3-Opus had a similar level of performance in this task, they both generated information of better quality than Elicit. As mentioned, Claude, rather than GPT, is chosen to extract information from the 319 papers to prevent potential data leakage.
4.2 Variant Generation
@@
After the experiment settings and conclusions are extracted, the next step is to generate variants based on the true conclusion, since the goal is to see if the LLM could select the true one under distraction. Inspired by Luo et al. Luo et al. (2024)’s prediction of neuroscience results by LLMs, for each conclusion from an experiment, the framework prompts the original conclusion and its two variants to GPT: its reversed variant and unrelated variant. As a result, the framework will make the prediction by choosing one from three options.
Figure 5: Variant Generation Process
@@
Specifically, as shown in Figure 5, the framework initially prompts the original conclusion to Claude, which follows the instructions to generate the reversed variant of the original conclusion. The reversed variant means that the direction of the conclusion is inverted. For example, if one conclusion is “receiving housing vouchers reduces quarterly employment rates,” its reversed variant is “receiving housing vouchers increases quarterly employment rates.” Next, the framework prompts both the original and reversed conclusions to Claude to generate the unrelated variant, which typically indicates that there is no correlation between entities of interest. Following the same example, the unrelated variant is “There is no relationship between receiving housing vouchers and quarterly employment rates.”
4.3 Prediction

### v2

@@ Abstract.
Large language models (LLMs) have demonstrated unprecedented emergent capabilities, including content generation, translation, and simulation of human behavior. Field experiments, on the other hand, are widely employed in social studies to examine real-world human behavior through carefully designed manipulations and treatments. However, field experiments are known to be expensive and time consuming. Therefore, an interesting question is whether and how LLMs can be utilized for field experiments. In this paper, we propose and evaluate an automated LLM-based framework to predict the outcomes of a field experiment. Applying this framework to 276 experiments about a wide range of human behaviors drawn from renowned economics literature yields a prediction accuracy of 78%. Moreover, we find that the distributions of the results are either bimodal or highly skewed. By investigating this abnormality further, we identify that field experiments related to complex social issues such as ethnicity, social norms, and ethical dilemmas can pose significant challenges to the prediction performance.
Large Language Models, Field Experiments
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: SciSoc LLM Workshop; August 6th, 2025; Torontoisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Collaborative and social computing systems and tools
1. Introduction
@@
Field experiments allow researchers to manipulate variables of interest in a real-world setting, such as human behaviors like ad clicking or donation giving, establishing causal relationships between interventions and outcomes. They typically begin by designing an intervention aligned with a specific research question, randomly assigning participants to treatment or control groups, and then measuring outcomes under natural conditions. The resulting data are collected and analyzed to evaluate the causal impact of the intervention (Levitt and List, 2009). It is adopted by a wide range of disciplines across academia and industry, such as finance (Berge et al., 2015; Du et al., 2020), marketing (Fradkin and Holtz, 2023; Li et al., 2020), and organizational studies (Hoogendoorn et al., 2017; Luciano et al., 2025). In recent years, online field experiments, also known as A/B testing, have revolutionized numerous online platform designs, advertising strategies, etc. While effective, field experiments are also known to be expensive and time consuming. For example, some experiments may take months or years to conduct (Minson et al., 2023). Moreover, recruiting high-quality participants for the experiments is challenging and costly (Anastasi et al., 2024), potentially affecting the experiment’s outcomes.
@@
In recent years, Large language models (LLMs) have demonstrated unprecedented emergent capabilities, including content generation, translation, and simulation of human behavior. For example, existing studies have explored the alignments between simulated data generated by LLMs and real data collected from human participants across various aspects, including human responses, traits (Dubois et al., 2024), moral standards (Dillion et al., 2023), preferences (Namikoshi et al., 2024), and emotions (Gao et al., 2023a). Therefore, it is becoming increasingly popular to leverage LLMs for simulating human behaviors (Xu et al., 2024; Gao et al., 2023b). For example, there are a handful of studies in which scholars have successfully instructed LLMs to replicate existing lab experiments across several disciplines, including psychology (Aher et al., 2023; Cui et al., 2024), sociology (Leng and Yuan, 2023; Manning et al., 2024; Hewitt et al., 2024), and economics (Aher et al., 2023; Horton, 2023). Their aim is to replicate existing lab experiments by treating LLMs as participants in lab experiments.
@@
However, directly applying these works to field experiment simulations in our context has several challenges. First, they mainly focused on lab experiment settings. But field experiments are inherently more challenging to conduct than laboratory experiments, due to diverse participant backgrounds, complex workflows, and multifaceted treatment designs. Second, most of the previous studies mainly relied on manual processes, and because of that, tested a handful of experiments with a limited scope of topics. While some recent studies tested on relatively large data, the selected experiments were limited to Likert-based psychological or social surveys (Cui et al., 2024; Hewitt et al., 2024). In order to explore LLMs’ capabilities, robustness, and generalizability in field experiment simulation, we need to test them on a much larger scale and broader range of experiments on different topics.
In this paper, we fill the literature gap by proposing an automated LLM framework to predict the outcome of a wide range of field experiments. Our framework has several major components, such as an information extraction module, a variant generation module, and a prediction module. Specifically, the information extraction module extracts key experiment settings, while the variant generation module generates false variants as distractors to confuse LLMs. Finally, the prediction modules leverage two prompt templates with a Chain-of-Thought design, prompting the LLM to predict the outcomes of a field experiment.
We test this framework on 276 field experiments reported in premier academic journals from 2000 to 2024. Those experiments contain a total of 1261 conclusions with a wide range of topics such as labor-market discrimination, educational incentives, household finance behavior, and the impact of healthcare enrollment. Without any alignment techniques or fine-tuning, our framework achieves an average prediction accuracy of 78%. We also test for data memorization effects by examining prediction results on recent experiments that appeared in 2024, which would be less likely to be included in the training data of LLMs. More interestingly, we also find that the prediction results are either bimodal or highly skewed. For example, our framework achieves nearly 100% prediction accuracy on 71% of conclusions while it completely fails to predict 18% of the conclusions with close to 0% accuracy. Further analyses reveal that our LLM-based framework has limitations in predicting experiments related to complex social issues such as ethnicity, social norms, and ethical dilemmas.
Our research makes several contributions:
  • 1
@@
    We extend the current literature on LLMs’ emergent capabilities by demonstrating that LLMs can simulate field experiments by predicting conclusions. While using LLMs to simulate human behavior is well studied in recent years, to the best of our knowledge, this is the first work in the literature to replicate large-scale field experiments that require a more complex environment setting and workflow design.
  • 2
@@
    Our proposed framework enables the prediction of field experiment conclusions in a fully automated, large-scale fashion. As a result, our framework is robust and can be generalized to a wider range of downstream applications in field experiments.
  • 3
    This paper examines the prediction performance and reveals the limitations of LLMs in field experiment simulations.
2. Literature Review
@@
While past literature mainly focused on agent-based social simulation (Van Heeswijk et al., 2020; Zhang et al., 2020; Chu and Shen, 2006), there is an increasing trend to adopt LLMs as simulation tools. Existing studies have found that LLMs’ ability to simulate human behavior stems from their possession of human-like reasoning skills and their adaptivity to personas of diverse characters (Shanahan, 2024; Taubenfeld et al., 2024a). Upon those features of LLMs, Aher et al. (Aher et al., 2023) proposed the concept of ”Turing Experiment”, in which LLMs are profiled as synthetic participants of experiments with integrated prompt of experimental settings and demographic information. Similarly, Horton (Horton, 2023) demonstrated LLMs’ ability to simulate lab experiments and promoted it as a method of experimental pilot testing. The usage of LLMs in Horton (Horton, 2023)’s work is similar to Aher et al. (Aher et al., 2023), consisting of two stages: prompting LLMs as synthetic experimental participants and collecting responses from conversations, reporting a few successful simulation cases of lab experiments. Leng & Yuan (Leng and Yuan, 2023) harnessed a three-phase procedure to complete the lab experimental simulation, which includes the initialization phase, interaction phase, and decision analysis phase. In the initialization phase, separate conversations of GPT-4 are prompted as vanilla experimental participants without specifying demographic information. Then, in the interaction stage, synthetic participants are prompted with the actions of other participants and asked to take actions and rationales according to the experiment design. Last, agents’ actions and rationales are collected for analysis to conclude. Leveraging this procedure, the study simulates five existing lab experiments. Manning et al. (Manning et al., 2024) proposed an automated framework for various lab experimental simulations. Although the entire workflow is divided into seven steps, it is essentially similar to the manual procedures of the three papers above. However, the sole input is the backstory of the experiment. Based on the input information and powered by an LLM, the framework continues with subsequent steps: identifying dependent and independent variables, generating treatment values, profiling agents, organizing the interaction of agents, collecting data, and establishing causal relationships and conclusions. The paper reports the results of four social scenarios. Besides the aforementioned studies, which were tested only on a small scale of cases, there are extant papers that implemented large-scale testing on LLM-based experimental replication. Cui et al. (Cui et al., 2024) replicated 154 psychological experiments. Specifically, they profiled LLMs as either students or adults, retrieving Likert-like responses from LLMs, which is oversimplified compared to most lab and field experiments. By comparison, Ashokkumar et al. (Hewitt et al., 2024) extended such a massive replication to 70 more complex survey experiments, though their selected experiments were limited exclusively to US social surveys and did not include any field experiments in the primary tests.
@@
In summary, while existing studies discussed above pioneered the application of Large Language Models (LLMs) to laboratory experiment simulations, several research gaps remain unaddressed. For instance, existing methods are not tested on field experiments that are inherently more complex than laboratory experiments, encompassing diverse participant backgrounds, more intricate workflows, and multifaceted treatment designs. Besides, despite extensive discussions on LLM bias (Chen et al., 2025; Wang et al., 2023; Du and Zhang, 2024) and the fact that recognized biases in LLMs—such as gender (Khorramrouz et al., 2023) and social norm bias (Ren et al., 2024)—can compromise performance on downstream tasks (Harris et al., 2022), only one of these studies mentioned the impact of certain topics on the replication successful rate of psychological surveys (Cui et al., 2024). It is still unclear how the joint effect of topics and sentiments would undermine the fidelity of field experiment prediction. Another limitation is that these studies tested their simulation strategies on only a small number of experiments, restricting generalizability. To address these gaps, we evaluated our proposed framework at scale on field experiments from published papers, demonstrating not only that LLMs can accurately predict the outcomes of established experiments, but also clarifying boundary conditions under which they cannot provide reliable experimental predictions. Furthermore, current automated simulation framework (Manning et al., 2024; Cui et al., 2024; Hewitt et al., 2024) are incompatible with scenarios involving complex treatments involving human-object interactions and limit only to lab experiments with Likert response or survey. By contrast, our framework offers a broadly adaptable approach that supports experiments across a wide range of contexts.
3. Data Collection and Filtering
@@
To explore LLMs’ ability to predict field experiments in a large-scale fashion, we first need to collect existing field experiments. Inspired by prior studies that focus on using LLMs to simulate human behaviors in lab experiments from existing psychology literature, we also consider field experiments studied in premier journals in economics, such as The Review of Economic Studies, American Economic Review, Journal of Political Economy, and The Quarterly Journal of Economics. It is worth noting that we focus on field experiments in economics because they are generally a larger scale in terms of participant size and robust in terms of careful design.
Figure 1. The Data Collection Workflow.
[…]
A flowchart illustrating the data collection process: from filtering 6544 papers to applying automated Claude-based checks and manual verification, resulting in 276 selected field experiments.
@@
Figure 1 shows the data collection and filtering workflow. Initially, 6544 papers containing keywords related to field experiments published between 2000 and 2024 of these top journals were selected. Then, we applied a two-layer verification process powered by Claude (Claude-3-opus-20240229). First, we prompted the title and abstract of each paper to Claude and asked it to judge if the paper designs and implements a field experiment. Upon this, the second verification prompted the entire paper to Claude, asking the same question. This strategy balances the accuracy of verification and the cost of calling Claude, as prompting the entire paper exponentially increases the cost. A final rule-based manual check to ensure the fidelity of the automated selection and filter out 276 papers for testing. More details about the manual check are available in Section 4.4. It is also worth mentioning that using Claude as the verification tool, instead of GPT, prevents potential data leakage as GPTs are leveraged as the prediction tool in our framework. The distribution of the selected papers is shown in Figure 7 (In Appendix A).
4. Framework
Figure 2. Prediction Framework.
[…]
A block diagram showing the three-stage LLM prediction framework: Extraction, Variant Generation, and Prediction, where Claude handles preprocessing and GPT makes final predictions.
@@ 4. Framework
We present the details of our automated framework for predicting field experiments in this section. Figure 2 shows the workflow. Overall, our framework is divided into three stages: Extraction from Papers, Variant Generation, and Prediction. Notably, Claude (Claude-3-opus-20240229) powers all preprocess tasks in the first two stages, whereas GPT completes the prediction at the last stage. We use two different LLMs in different stages to prevent potential data leakage from one another.
4.1. Extraction
@@
Specifically, the framework uses Claude to extract information related to a field experiment from the selected paper. To realize this, the framework leverages a manually crafted prefixed prompt, which has proven to be efficient for various downstream tasks (Brown et al., 2020). As shown in Figure 8 (In Appendix B.1), the prompt template contains a placeholder ”${Paper}$”, an information form consisting of bullet points from ”A” to ”G”, and clear instructions that ask the LLM to extract information according to the form from the paper. As underscored in the prompt, the first six bullet points ”A” to ”F” are key experimental settings that shape the experiment context, whereas the last point ”G” is about conclusions that are true outcomes in the prediction task.
@@
Based on the response from Claude, the framework formulates ”Experiment Settings” directly from bullet points ”A” to ”F”, while it polishes point ”G” to generate ”Experiment Conclusions”. Specifically, the raw response regarding point ”G” is a paragraph containing multiple conclusions of the field experiment. To separate that paragraph into standalone conclusions, the framework calls a new Claude session, prompting the raw paragraph and related instructions to it, finally getting ”Experiment Conclusions” in return. Breaking complex tasks into subtasks improves the performance of LLM-driven workflows (Zhang and Gao, 2023), which is the main reason for completing the extraction and separation of conclusions in different Claude sessions.
4.2. Variant Generation
@@
After the experiment settings and conclusions are extracted, the next step is to generate variants based on the true conclusion, since the goal is to see if the LLM could select the true one under distraction. Inspired by Luo et al. (Luo et al., 2024)’s prediction of neuroscience results by LLMs, for each conclusion from an experiment, our framework prompts the original conclusion and its two variants to GPT: a reversed variant and an unrelated variant. As a result, the framework will make the prediction by choosing one of three options.
@@
Specifically, as shown in Figure 9 (In Appendix B.2), the framework initially prompts the original conclusion to Claude, which follows the instructions to generate the reversed variant of the original conclusion. The reversed variant means that the direction of the conclusion is inverted. For example, if one conclusion is ”receiving housing vouchers reduces quarterly employment rates,” its reversed variant will be ”receiving housing vouchers increases quarterly employment rates.” Next, the framework prompts both the original and reversed conclusions to Claude to generate the unrelated variant, which typically indicates that there is no correlation between entities of interest. Following the same example, the unrelated variant will be ”There is no relationship between receiving housing vouchers and quarterly employment rates.”
4.3. Prediction